1.) Put the conic in standord form by completing the square, please identify which conic it is, and then graph the conic section. Also label each relevant point or asymptotes. Sketch your x - y axis and the conic, and labl the scale on the axes.

c. -9x2 + y2 – 72x – 153 = 0

example sketches:

Ex. 1 Sketch (x-5) V =2 center: (5,2) radius: 2

2.)Center , horizontal: 2 units left/right 3 units up/down (1,3) 13,0) (1,-3)
Answer #1

1) -92+y^-72% -153 = 0 = -97-72x – 124 ty_920 0-9 (78x+16) +4–920 - 96*747-y-9 f (x+4) - Tart 2 Cathy - 2 = 1 or, - Comis X =